# Architecting mechanisms of damage in topological metamaterials

### v1

@@ Conclusions
Methods
In the simulations, we represent the lattice as a mass-spring network with both axial and torsional springs providing NN and NNN interactions. The energy in this system is defined by
where: $N_{\mathrm{b}}$ and $N_{\mathrm{r}}$ are the number of axial and torsional springs in the system; $k_{\mathrm{i}}$ and $c_{\mathrm{s}}$ are the axial and torsional spring stiffness; $l_{\mathrm{i}}$ and $l_{\mathrm{i}}^{\mathrm{0}}$ are the current length and original length of the axial spring; and, $\alpha_{\mathrm{s}}$ and $\alpha_{\mathrm{s}}^{\mathrm{0}}$ are the current angle and original angle between bonds $i$ and $l$. Notice that this formulation captures large deformations. $l_{\mathrm{i}}$ and $\alpha_{\mathrm{s}}$ can be written in terms of the site degrees of freedom $x_{\mathrm{j}}$ and, within each applied displacement step $\delta$, we minimises the force at each site ($\partial E/\partial x_{\mathrm{j}}$) below a specified tolerance ($F_{\mathrm{tol}}$) using the FIRE 2.0 algorithm [35, 36].
To accurately capture the bending of the bonds, we further discretise the bonds into three elements. This level of discretisation was found to sufficiently capture the response of the lattice for the SR and material properties simulated.
An element has reached failure if
where $\sigma_{\mathrm{A}}$ and $\sigma_{\mathrm{B}}$ are the axial and bending stresses in the element, and $\sigma_{\mathrm{A}}^{\mathrm{u}}$ and $\sigma_{\mathrm{B}}^{\mathrm{u}}$ are the material failure stress values in axial and bending respectively. While only the material failure is included explicitly, buckling related failure modes are captured by this model. When a bond becomes unstable, further rotation is possible without significantly increasing the applied force [37]. This far-from-threshold rotation will eventually result in the failure of an element, which is simulated by splitting the site associated with the failed element into two, creating a new edge within the lattice. Further information about the method can be found in SI Appendix.
The material properties are taken from that of a relatively rigid, but brittle, material [38], with $E=10$GPa, $\sigma_{\mathrm{A}}^{\mathrm{u}}=65$MPa and $\sigma_{\mathrm{B}}^{\mathrm{u}}=126$MPa. The maximum allowable residual force ($F_{\mathrm{tol}}$) was set to 0.0025N in the simulations following a sensitivity study.
Acknowledgements.
@@ Appendix A Method
Figure 7: (a) NN and NNN bonds. (b) Discretisation of the lattice. Splitting of sites with: (c) two NN bonds; (d) three NN bonds; and, (e) four NN bonds.
Spring stiffness
@@
Site Splitting
When either side of the bond meets the failure criteria (Eq. (7) within the main text), we separate the bond(s) by splitting the site into two and reassigning the bonds to this new site. This is done arbitrarily when two elements are connected to a site. The associated torsional spring stiffness ($c_{\mathrm{s}}$) is also set to 0. In Fig. 7c we show this process, where a yellow circle represents the new site. When three or four elements are connected at a site we choose to disconnect the elements about the two torsional springs whose angle changes the most i.e. if two elements are rotating together about the site they will separate from the other element(s) together (Fig. 7d,e). $c_{\mathrm{s}}$ of the two torsional springs associated with the largest rotations is then also set to 0. If, in this step, one element is separated from three elements, a new torsional spring is added to the set of three elements, connecting the adjacent elements that were previously on either side of the single separated element (Fig. 7e). No torsional springs are added to the system if only two elements remain after splitting the site (Fig. 7d,e).
Algorithm
@@ Appendix B Evanescent Mode Visualizations
In this section, we explain the methods used to visualize the form of evanescent modes associated with bulk discontinuities. It is important to mention that although the solutions we visualize given a surface disturbance are exact, the (functional) form of this disturbance has to be prescribed following certain assumptions. The visualizations are thus an instructive means of designing the lattices. For instance, in the case of FM on the crack surface of LP lattices, we use the qualitative assumption of a strongly localised disturbance on the crack tip, which can be motivated through fracture mechanics on micropolar media [41], as well as our simulation results (see Fig.2i). For RP lattices, which do not possess a significant spike, we have instead used a sigmoid prescription, which aims to emulate the relaxation of the lattice on the crack-edge, and the fact that the disturbance approaches a uniform actuation in the bulk. Finally, although our examples are focused on FM visualisations, note that the same method can be applied for SSS.
Figure 8: Demonstrations of disturbances, (a-b) without and (c) with boundary conditions, for LP1 - all continuum plots depict the modulus $|\phi|$, of the otherwise complex FM field. In all cases, green solid lines represent free edges, with (a): a distortion on the edge made up of a single wavenumber $q_{\mathrm{1}}=2\pi s/N$, i.e. $\tilde{\phi}_{m}=\delta_{ms}$, with a visualization of the distorted lattice fading into the continuum and (b,c): Gaussian distortions of varying widths. The dashed black line in (c) represents a boundary with rollers, delineating the top half domain, which is virtual and contains the mirror image of the field in the lower half - note the two vortexes visualising the mirroring.
@@ B.1 Bulk-Boundary Correspondence
As shown in [33], in the limit of linear elasticity, the existence of uniform strain, zero energy (e.g. GH) modes imply the existence of spatially modulated versions of the same strain that also cost no energy. For idealized, zero bending stiffness kagome lattices, the zero energy distortion on the unit cell can be captured by a single parameter $\phi$ [39, 40], which, when allowed to vary spatially defines a type of micro-deformation field $\phi(\bm{x})$. In the linear elastic limit, geometric compatibility has to be enforced, which constrains the functional form for $\phi(\bm{x})$, effectively reducing the degrees of freedom of the function from two independent coordinates $\phi\equiv\phi(x,x)$ into a combination of the two $\phi\equiv\phi(x+\lambda x)$.
This phenomenon, known as bulk-boundary correspondence, can be seen for the full-field solution because the energy cost of the deformation must vanish. In Fourier Space we can relate the orthogonal components of the wavenumbers describing the (floppy) distortion of the lattice, $q_{\mathrm{1}}$ and $q_{\mathrm{2}}$ through this constraint:
@@
allowing us to write, e.g. $q_{\mathrm{2}}\equiv G(q_{\mathrm{1}})$, prescribing one wavenumber component in terms of the other. Given that $C(q_{\mathrm{1}},q_{\mathrm{2}})$ contains all information about the lattice, this model does not make any homogenizing assumptions - $q_{\mathrm{2}}=G(q_{\mathrm{1}})$ is exact, $\forall q_{\mathrm{1}}\in\left[-\pi/a,\pi/a\right]$. Note that, as $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$ although we focus on FM, the extension to the discussion of SSS is straightforward. Now, using this result, consider the deformation field $\phi(\bm{x})$, written as an inverse Fourier Transform:
@@
with $l,m$ indexing the $x,y$ locations in the grid, and where we are summing over the wavenumbers $q_{\mathrm{1}}\equiv\frac{2\pi n}{N}$ $G(q_{\mathrm{1}})$ being the geometry dependent function we retrieve through the zero energy constraint (12). Then, the Fourier components $\tilde{\phi}_{n}$ can be found through a prescription of the field on N points e.g. on the horizontal line defined by $m=m_{0}$,
@@
Boundary Conditions
A FM, in general, distorts the lattice in a way that does not respect the Boundary Conditions (BC). An equivalent way of viewing the continuation of the disturbance around the boundary constraint is to consider a secondary disturbance $\phi_{b}(x,y)$ localized on the boundary, which, when superimposed with the primary disturbance localized on the crack tip $\phi_{0}(x,y)$, has to give a displacement field that respects the BCs. For the strongly localised crack tip actuation, there are two sets of Dirichlet BCs we use in this work: (1) rollers ($\delta y=0$) and (2) pinned ($\delta x=0$, $\delta y=0$). Although pinning the sites lifts the FM [20], the roller boundary conditions may allow for their existence by a reflection-like propagation that we examine below.
For y-mirror-symmetric unit cells (LP, RP, TK), roller boundary conditions can be satisfied by superimposing two images of the disturbance, mirrored around the boundary. This is because the mirroring achieves the desired $\delta y_{b}=-\delta y_{0}$, on the boundary. Furthermore, the continuation of the real domain into its virtual counterpart is compatible with the unit cell geometry, as the unit cell is chiral. For achiral unit cells that still belong in the equivalence classes (±,±,±)plus-or-minusplus-or-minusplus-or-minus($\pm$,$\pm$,$\pm$)( ± , ± , ± ) or (,±,±)minus-or-plusplus-or-minusplus-or-minus($\mp$,$\pm$,$\pm$)( ∓ , ± , ± ), mirroring will cause a discontinuity in the lattice tiling around the boundary. Nonetheless, mirroring would still satisfy the BC, with the extension of the FM from one domain to the other refracting the disturbance. Numerically, the FM can be constructed together with its reflection by obtaining the primary amplitude on the boundary $\phi_{0}(x,y_{b})$, and then using this as the new prescription for the secondary mode $\phi_{b}(x,y)$.
For lattices in the (±,,±)plus-or-minusminus-or-plusplus-or-minus($\pm$,$\mp$,$\pm$)( ± , ∓ , ± ) or (±,±,)plus-or-minusplus-or-minusminus-or-plus($\pm$,$\pm$,$\mp$)( ± , ± , ∓ ) equivalence classes and initial geometries with a horizontal crack, like the ones considered in this work, both modes localize on the same horizontal edge, and the disturbance cannot be canceled on the boundary by a mode localizing there. It is not directly clear what combination of modes localizing on the remaining edges is induced in the lattice, so as to respect the boundary conditions. As a result, we limit our visualizations to the case of open boundary conditions for these modes.
@@ Appendix C Behaviour of additional topologically distinct lattices
Dominant lattice behaviour
@@
Figure 9: Snapshot of: (a) axial energy portion in the entire lattice; and, (b) axial stress fraction in the element that fails preceding the first failure event. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for RTK at: (c) SR = 3; and, (d) 30. The difference in angle is visualised through the colour directly to the right of each unit cell. Damage path for: (e) RTK; and, (f) RDP1. The damage process is visualised with a line traversing the domain. The line connects separated nodes, where fracture occured, within the lattice. For clarity, we do not show the micro-structure throughout the domain. The damage process for SR = 3, 10, 30 is plotted with green, black and blue lines and associated rectangular markers respectively (refer to (b)). Arrows indicate the direction of damage. Normalised maximum bending stress in cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (g,i) RDP1 (0.1,-0.1,0.1); and, (h,j) RDP3 (0.16,-0.04,0.16) at SR = 30. (k) $\theta$ spectrum for RDP1 and RDP3. (l) Damage path for RDP1 and RDP3 at SR = 30.
Dominant lattice failure
@@ Appendix D Additional constraints
Dominant SSS
Figure 10: Size effect study for LP1 with AR = 0.14: (a) $\kappa^{\mathrm{SSS}}$; (b) $\theta$; (a) axial energy portion vs length of domain at $\epsilon_{\mathrm{n}}=0.01$; and, total normalised stress profile ($\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$) for lattice with: (d,h) n = 8, m = 50; (e,i) n = 16, m = 100; (f,j) n = 32, m = 200; and, (g,k) n = 48, m = 300.
@@
Failure evolution for LP2 lattice with larger length
@@
Stress profile in lattices from (--,$+$,--) equivalence class
@@
Figure 12: Failure path, normalised axial and normalised bending stress for: (a-c) LUP1; (d-f) LUP2; and, (g-i) LUP3 respectively at SR = 30 with pinned boundary conditions. In (a), (d) and (g), we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows on either side of these zones. If the SSS can’t develop we only plot the direction of the SSS with a single arrow.
Appendix E Asymmetric lattices from ($\mp$,$\pm$,$\pm$) equivalence class
All lattices reported in the main text from the ($\mp$,$\pm$,$\pm$) equivalence classes had symmetric FM and SSS forms. Here, we provide additional simulations for asymmetric forms. These additional lattices are: LP4 (0.1, -0.05, -0.15); RP4 (-0.1, 0.05, 0.15); LP5 (0.04, -0.25, -0.15); and, LP6 (0.04, -0.25, -0.05). All simulations are for SR = 30.
@@ Appendix E Asymmetric lattices from ($\mp$,$\pm$,$\pm$) equivalence class
Figure 13: Bending stress in: (a) LP4; and, (b) RP4. SSS development at the long-wavelength limit and normalised stress in domain for: (c, e) LP5; and, (d, f) LP5. All simulation results are for SR = 30 with pinned boundary conditions. (g) $\theta$ spectrum for LP4, LP5 and LP6.

### v2

@@ 6 Conclusions
Acknowledgments
Funding:
@@ Acknowledgments
Author contributions:
L.d.W. and M.A.D. designed the research; M.A.D. supervised and acquired funding; L.d.W. developed the numerical analysis; L.d.W., M.C., and M.A.D. performed research; L.d.W., M.C., and M.A.D. analyzed data; L.d.W. wrote original draft; and L.d.W., M.C., and M.A.D. wrote and edited the paper.
Competing interests:
There are no competing interests to declare.
@@ A.1 Discretisation
A.2 Spring stiffness
@@ A.3 Stress in bonds
A.4 Algorithm
@@ Appendix B Evanescent Mode Visualizations
In this section, we explain the methods used to visualize the form of evanescent modes associated with bulk discontinuities. It is important to mention that although the solutions we visualize given a surface disturbance are exact, the (functional) form of this disturbance has to be prescribed following certain assumptions. The visualizations are thus an instructive means of designing the lattices. For instance, in the case of FM on the crack surface of LP lattices, we use the qualitative assumption of a strongly localised disturbance on the crack tip, which can be motivated through fracture mechanics on micropolar media (Berkache et al., 2022), as well as our simulation results (see Fig.2i). For RP lattices, which do not possess a significant spike, we have instead used a sigmoid prescription, which aims to emulate the relaxation of the lattice on the crack-edge, and the fact that the disturbance approaches a uniform actuation in the bulk. Finally, although our examples are focused on FM visualisations, note that the same method can be applied for SSS.
@@ B.1 Bulk-Boundary Correspondence
In the limit of linear elasticity, the existence of uniform strain, zero energy (e.g. GH) modes imply the existence of spatially modulated versions of the same strain that also cost no energy (Rocklin et al., 2017). For idealized, zero bending stiffness kagome lattices, the zero energy distortion on the unit cell can be captured by a single parameter $\phi$ (Sun et al., 2012; Nassar et al., 2020), which, when allowed to vary spatially defines a type of micro-deformation field $\phi(\boldsymbol{x})$. In the linear elastic limit, geometric compatibility has to be enforced, which constrains the functional form for $\phi(\boldsymbol{x})$, effectively reducing the degrees of freedom of the function from two independent coordinates $\phi\equiv\phi(x,x)$ into a combination of the two $\phi\equiv\phi(x+\lambda x)$.
This phenomenon, known as bulk-boundary correspondence, is also apparent in the full-field (non-linear) solution, where it can be understood as a restatement of the zero energy requirement. In Fourier space we can relate the orthogonal components of the wavenumbers describing the (floppy) distortion of the lattice, $q_{\mathrm{1}}$ and $q_{\mathrm{2}}$ through this constraint:
@@
allowing us to write, e.g. $q_{\mathrm{2}}\equiv G(q_{\mathrm{1}})$, prescribing one wavenumber component in terms of the other. Given that $C(q_{\mathrm{1}},q_{\mathrm{2}})$ contains all information about the lattice, this model does not make any homogenising assumptions - $q_{\mathrm{2}}=G(q_{\mathrm{1}})$ is exact, $\forall q_{\mathrm{1}}\in\left[-\pi/a,\pi/a\right]$. Note that, as $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$—although we focus on FM, the extension to the discussion of SSS is straightforward. Now, using this result, consider the deformation field $\phi(\boldsymbol{x})$, written as an inverse Fourier Transform:
@@
with $l,m$ indexing the $x,y$ locations in the grid, and where we are summing over the wavenumbers $q_{\mathrm{1}}\equiv 2\pi n/N$; $G(q_{\mathrm{1}})$ being the geometry dependent function we retrieve through the zero energy constraint (13). Then, the Fourier components $\tilde{\phi}_{n}$ can be found through a prescription of the field on N points e.g. on the horizontal line defined by $m=m_{0}$,
@@ B.2 FM Prescriptions
B.3 Boundary Conditions
A FM, in general, distorts the lattice in a way that does not respect the Boundary Conditions (BC). An equivalent way of viewing the continuation of the disturbance around the boundary constraint is to consider a secondary disturbance $\phi_{b}(x,y)$ localized on the boundary, which, when superimposed with the primary disturbance localized on the crack tip $\phi_{0}(x,y)$, has to give a displacement field that respects the BCs. For the strongly localised crack tip actuation, there are two sets of Dirichlet BCs we use in this work: (1) rollers ($\delta y=0$) and (2) pinned ($\delta x=0$, $\delta y=0$). Although pinning the sites lifts the FM (Kane and Lubensky, 2014), the roller boundary conditions may allow for their existence by a reflection-like propagation that we examine below.
Figure 10: Continuum surface mode visualizations. Demonstrations of disturbances with, (a) open and (b) roller boundary conditions, for LP1. These continuum plots depict the modulus $|\phi|$, of the otherwise complex FM field. In all cases, green solid lines represent free edges, with Gaussian distortions of varying widths. The dashed black line in (b) represents a boundary with rollers, delineating the top half domain, which is virtual and contains the mirror image of the field in the lower half - note the two vortexes highlighting the mirroring.
For y-mirror-symmetric unit cells (LP, RP, TK), roller boundary conditions can be satisfied by superimposing two images of the disturbance, mirrored around the boundary. This is because the mirroring achieves the desired $\delta y_{b}=-\delta y_{0}$, on the boundary (see Fig. 10b). Furthermore, the continuation of the real domain into its virtual counterpart is compatible with the unit cell geometry, as the unit cell is chiral. For achiral unit cells that still belong in the equivalence classes (±,±,±)plus-or-minusplus-or-minusplus-or-minus($\pm$,$\pm$,$\pm$)( ± , ± , ± ) or (,±,±)minus-or-plusplus-or-minusplus-or-minus($\mp$,$\pm$,$\pm$)( ∓ , ± , ± ), mirroring will cause a discontinuity in the lattice tiling around the boundary. Nonetheless, mirroring would still satisfy the BC, with the extension of the FM from one domain to the other refracting the disturbance. Numerically, the FM can be constructed together with its reflection by obtaining the primary amplitude on the boundary $\phi_{0}(x,y_{b})$, and then using this as the new prescription for the secondary mode $\phi_{b}(x,y)$.
For lattices in the (±,,±)plus-or-minusminus-or-plusplus-or-minus($\pm$,$\mp$,$\pm$)( ± , ∓ , ± ) or (±,±,)plus-or-minusplus-or-minusminus-or-plus($\pm$,$\pm$,$\mp$)( ± , ± , ∓ ) equivalence classes and initial geometries with a horizontal crack, like the ones considered in this work, both modes localize on the same horizontal edge, and the disturbance cannot be canceled on the boundary by a mode localizing there. It is not directly clear what combination of modes localizing on the remaining edges is induced in the lattice, so as to respect the boundary conditions. As a result, we limit our visualizations to the case of open boundary conditions (e.g. Fig. 10a) for these modes.
@@ Appendix C Behaviour of additional topologically distinct lattices
C.1 Dominant lattice behaviour
@@
Figure 11: Dominant behaviour and failure of additional lattices. Snapshot of: (a) axial energy portion in the entire lattice; and, (b) axial stress fraction in the element that fails preceding the first failure event. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for RTK at: (c) SR = 3; and, (d) 30. The difference in angle is visualised through the colour directly to the right of each unit cell. Damage path for: (e) RTK; and, (f) RDP1. The damage process is visualised with a line traversing the domain. The line connects separated nodes, where fracture occured, within the lattice. For clarity, we do not show the micro-structure throughout the domain. The damage process for SR = 3, 10, 30 is plotted with green, black and blue lines and associated rectangular markers respectively (refer to (b)). Arrows indicate the direction of damage. Normalised maximum bending stress in cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (g, i) RDP1 (0.1,-0.1,0.1); and, (h, j) RDP3 (0.16,-0.04,0.16) at SR = 30. (k) $\theta$ spectrum for RDP1 and RDP3. (l) Damage path for RDP1 and RDP3 at SR = 30.
C.2 Dominant lattice failure
@@ Appendix D Additional constraints
D.1 Dominant SSS
@@
Figure 12: Size effect study for LP1 with AR = 0.14. (a) $\kappa^{\mathrm{SSS}}$; (b) $\theta$; (C) axial energy portion vs length of domain at $\epsilon_{\mathrm{n}}=0.01$; and, total normalised stress profile ($\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$) for lattice with: (D, H) n = 8, m = 50; (E, I) n = 16, m = 100; (F, J) n = 32, m = 200; and, (G, K) n = 48, m = 300.
@@
D.2 Failure evolution for LP2 lattice with larger length
@@
D.3 Stress profile in lattices from (--,$+$,--) equivalence class
@@
Figure 14: Stress development in LUP lattices with pinned boundary conditions. Failure path, normalised axial and normalised bending stress for: (a-c) LUP1; (d-f) LUP2; and, (g-i) LUP3 respectively at SR = 30 with pinned boundary conditions. In (a), (d) and (g), we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows on either side of these zones. If the SSS can’t develop we only plot the direction of the SSS with a single arrow.
Appendix E Achiral lattices from the ($\mp$,$\pm$,$\pm$) equivalence classes
All lattices reported in the main text from the ($\mp$,$\pm$,$\pm$) equivalence classes had FM and SSS with forms that where mirror symmetric wrt. the plane of the crack. Here, we provide additional simulations for lattices with asymmetric forms for these mode, i.e. achiral unit cell geometries. In particular we choose to visualise: LP4 (0.1, -0.05, -0.15); RP4 (-0.1, 0.05, 0.15); LP5 (0.04, -0.25, -0.15); and, LP6 (0.04, -0.25, -0.05). All simulations are for SR = 30.
@@ Appendix E Achiral lattices from the ($\mp$,$\pm$,$\pm$) equivalence classes
Figure 15: LP and RP lattice with achiral FM/SSS forms. Bending stress in: (a) LP4; and, (b) RP4. SSS development at the long-wavelength limit and normalised stress in domain for: (c, e) LP5; and, (d, f) LP5. All simulation results are for SR = 30 with pinned boundary conditions. (g) $\theta$ spectrum for LP4, LP5 and LP6.